OpenStudy (anonymous):

Medal and Fan for whoever answers A painting company charges $250 base plus $16 per hour. Another painting company charger $210 base plus $18 per hour. How long is a job for which the two companies costs are the same?

OpenStudy (anonymous):

Just solve the equation \[250+16x=210+18x\] for \[x\]

OpenStudy (akashdeepdeb):

Let 't' be the number of hours. Company A : 250 + 16t Company B : 210 + 18t Now, for the charges to be same, equate them! 250 + 16t = 210 + 18t 2t = 40 t = 20 hours Getting this? :D

OpenStudy (anonymous):

Yep thanks @AkashdeepDeb